Can someone please help with this question:
A rectangular piece of sheet metal has a length that is 4 in. less than twice the width. A square piece 2 in. on a side is cut from each corner. The sides are then turned up to form an uncovered box of volume 256 in^3. Find the length and width of the original piece of metal.
the answer I cam up with was way off.

Answer by rothauserc(4718)

length (l) = 2*w -4
Volume (V) = (2*w -4)*w*2 = 256
256 = 4w^2 -8w
divide both sides of = by 4
64 = w^2 -2w
put equation in standard form
w^2 -2w -64 = 0
use quadratic formula to solve this
w = (2 + square root((-2)^2 - 4*1*(-64))) / 2 = 9.06
w = (2 - square root((-2)^2 - 4*1*(-64))) / 2 = -7.06
therefore w = 9.06 and l = 14.12
check the answer with the V equation
14.12 * 9.06 * 2 = 256
255.9 approx 256
answer checks